# Do Large Language Model Understand Multi-Intent Spoken Language ?

### v2

@@ 5 Further Exploration
This section delves into a comprehensive analysis of the role and potential of LLMs in enhancing multi-intent SLU. Significantly, for cost-efficiency and replicability purposes, the experiments discussed within this section are uniformly conducted using 7B models.
5.1 Improvement Analysis
Large Language Models (LLMs) consistently surpass baseline models in performance across a comprehensive range of metrics within both the MixATIS and MixSNIPS datasets. To discern the performance differentials among the models, a methodical grouping was employed, predicated on the number of intents within MixATIS’s multi-intent dataset as captured in Table 2. It becomes evident that the escalation in the number of intents proportionately augments the difficulty for all models, as manifested in the diminished metrics across the board.
In a single-intent environment, the Vicuña-7B-v1.5 model demonstrates superior performance over the CLID(Roberta) model, achieving an 80.4% overall accuracy rate and marking an enhancement of 2.8%. As we venture into scenarios with two and three intents, a ubiquitous downward trend in model efficacy is noted. Yet, within these more demanding contexts, the Mistral-7B-Instruct-v0.1 model emerges as exceptional. It surpasses the CLID(Roberta) baseline by 6.5% in two-intent settings, achieving a 55.3% overall accuracy, the highest among the models under review. This distinction is sustained with three intents, where Mistral-7B-Instruct-v0.1 model attains an overall accuracy of 31.5%, topping the charts and revealing a 2.5% improvement over CLID(Roberta). The data suggests that Mistral-7B-Instruct-v0.1 possesses a significant aptitude for adeptly managing the complexities associated with multi-intent scenarios.
5.2 Sub Intent Instruction’s Efficacy in Multi-Domain SLU
Acknowledging that Large Language Models are extensively trained within diverse domains, our investigation extends to the Sub Intent Instruction (SII) and its role in bolstering LLMs’ proficiency for discerning and segmenting multi-intent utterances into discrete sub-intent components within the multi-domain MixSNIPS dataset. This enhancement ostensibly refines the overall multi-intent Spoken Language Understanding (SLU) capability—a claim substantiated by the data in Table 3.
Upon the integration of SII, although it does not improve significantly on Intent(Acc) and Slot(f1), a substantial increase in Overall Accuracy has been observed across all models, underlining SII’s pivotal role in boosting joint intent and slot recognition efficacy. For instance, Vicuna-7B-V1.5’s Overall(Acc) witnessed a significant rise of 3.5%, climbing from 78.9% to 82.4%. A similar enhancement was seen with Llama-2-7B-Chat, which recorded a robust improvement of 4.6%, reaching an Overall(Acc) of 83.5%. Mistral-7B-Instruct-v0.1 also experienced a notable gain of 4.1% in Overall(Acc), achieving 83.9%.
These pronounced increments in Overall(Acc) with the integration of SII emphasize its pivotal significance in ameliorating multi-intent SLU performance, necessitating accurate identification of the intricate links between various intents and their corresponding slots.
Model Slot(F1) Intent(Acc) Overall(Acc)
Vicuna-7B-v1.5 95.7 97.6 78.9
+SII 96.2 (↑0.5) 97.1 (↓0.5) 82.4 (↑3.5)
Llama-2-7B-chat 95.7 96.9 78.9
+SII 96.5 (↑0.8) 97.6 (↑0.7) 83.5 (↑4.6)
Mistral-7B-Instruct-v0.1 95.6 97.6 79.8
+SII 96.7 (↑1.1) 97.5 (↓0.1) 83.9 (↑4.1)
Table 3: Further experiments on the MixSNIPS dataset showing the impact of the sub intent instruction (SII). The brackets indicate the change in performance metrics with the SII approach compared to the original model.
@@ 5.2 Sub Intent Instruction’s Efficacy in Multi-Domain SLU
Figure 3: Performance comparison of models on MixATIS and MixSNIPS datasets. The figures depict Semantic Accuracy, Slot F1, and Intent Accuracy across different data ratios for Mistral, Vicuna, and Llama models (7B).
5.3 Correlation between Training Data Volume and Model Efficacy
@@
As illustrated in Figure 3, we probes the effect of varying volumes of training data on model proficiency, with a focus on three critical performance indicators: Semantic Accuracy, Slot F1 Score, and Intent Accuracy. Through progressive modulation of training data ratios at intervals of 0.2, 0.4, 0.6, 0.8, and 1.0. There appears to be a consistent and positive correlation between the data ratio and the amelioration of performance metrics across both the MixATIS and MixSNIPS datasets. This trend substantiates the hypothesis that an increase in dataset size can significantly bolster the model’s training efficacy, thereby refining accuracy and broadening the model’s generalization abilities. Upon closer inspection, the MixATIS and MixSNIPS datasets indicate that the Mistral-7B-Instruct-v0.1 model, starting at a less optimal performance with a data ratio of 0.2, demonstrates notable advancements as the dataset size increases. For instance, within the MixSNIPS dataset, there is an impressive ascent in Semantic Accuracy from 51.16% to 79.76%, Slot F1 from 76.18% to 95.64%, and Intent Accuracy from 72.99% to 97.64%. These substantial improvements underscore the model’s ability to enhance its learning and semantic processing capabilities with access to a larger corpus of data. Meanwhile, in the MixATIS dataset, Mistral-7B-Instruct-v0.1 achieves promising results in Semantic Accuracy and Slot F1. However, it appears to have compromised some aspect of Intent Accuracy in comparison to its counterparts, Llama2 and Vicuna.
6 Constructing Metrics for Evaluating LLMs in SLU
@@
To enhance the evaluation of LLMs in the domain of multi-intent SLU, we propose two novel metrics for a more nuanced analysis: Entity Slot Accuracy (ESA) and Combined Semantic Accuracy (CSA).
Entity Slot Accuracy (ESA): ESA measures the model’s ability to precisely identify entity slots within utterances, which are crucial for understanding the specific semantic elements of the spoken input.
Combined Semantic Accuracy (CSA): CSA amalgamates intent accuracy with ESA to yield a comprehensive gauge of a model’s capability in discerning the overall semantic framework of utterances. The metric is mathematically represented as follows:
$CSA=\text{Intent Accuracy}\times\text{ESA}$ (4)
Upon analyzing the performance of various Language Learning Models (LLMs) on the MixATIS and MixSNIPS datasets, the data reveals certain performance trends. The Mistral-7B-Instruct-v0.1 model outshines others in the MixATIS dataset with an ESA of 60.6%, and a CSA of 48.8%. In contrast, the Llama-2-13B model demonstrates a remarkable aptitude in the MixSNIPS dataset, achieving the highest ESA at 83.6% and the highest CSA at 81.7%. These scores reflect the model’s proficiency in handling the complexity of the dataset’s queries. It is worth noting that the ESA and CSA metrics provide a refined assessment of model performance by exclusively considering the value of entity slots, without taking into account the "O" tag typically used to denote tokens that do not correspond to any entity slot in the dataset’s utterances. This focus results in a more rigorous evaluation of an LLM’s ability to recognize and accurately predict the relevant information within a given input.
@@
Vicuna-7B-v1.5 53.3 43.5 79.1 77.2
Llama-2-7B-chat 53.1 43.8 79.3 76.8
Vicuna-13B-v1.5 56.2 46.9 80.9 78.9
Llama-2-13B 59.5 48.3 83.6 81.7
Mistral-7B-Instruct-v0.1 60.6 48.8 80.2 78.3
@@ 7 Conclusion
In this study, we have integrated LLMs into the multi-intent SLU framework, marking a leap forward in four critical areas: (1) In terms of modeling, we pioneered the application of LLMs for multi-intent SLU, offering fresh insights and systematically tackling the inherent complexities of the domain. (2) From a data perspective, we crafted entity slots optimized for LLMs within the multi-intent SLU context and established sub-intent instructions to probe the impact of sub-sentence components. The datasets refined for this purpose are named LM-MixATIS and LM-MixSNIPS. (3) Regarding performance, our investigation reveals that the LLM-based approach demonstrates impressive competitiveness, and in many instances, surpasses the capabilities of the existing state-of-the-art models in various benchmarks. (4) In the domain of metrics, we have introduced two novel measures: Entity Slot Accuracy (ESA) and Combined Semantic Accuracy (CSA), to facilitate a comprehensive and nuanced assessment of LLMs within multi-intent SLU. Through these advancements, we aspire to catalyze greater interest and exploration into the potential of LLMs to enhance multi-intent SLU systems.
@@ 8 Limitations
In this paper, we openly acknowledge several limitations of our approach: (1) Quantization Impact on Performance: Our implementation leverages QLoRA quantization to enhance training efficiency and reduce memory consumption. However, this approach might not reach the full efficacy of complete-parameter fine-tuning. The inherent trade-off in quantization is efficiency at the potential expense of some performance metrics. While it streamlines the process, it may inadvertently overlook subtle nuances that full-parameter models can capture. (2) Anticipated Enhancements through Data Selection and Prompt Refinement: Our current study does not encompass the advanced techniques of meticulous data selection or sophisticated prompt engineering. Recognizing this as a limitation, our future work will aim to incorporate these critical strategies. By deliberately curating datasets and enhancing prompt design in subsequent research, we intend to further refine the model’s performance, especially in areas where quantization may have compromised subtleties in linguistic understanding. These anticipated improvements will strive to strike a balance between efficiency and the depth of language comprehension, presenting a more comprehensive solution within the domain of Large Language Models.
Appendix A Related Work
A.1 Intent Detection and Slot Filling
Intent detection and slot filling are often interrelated, giving rise to the development of integrated models that facilitate interaction between intent and slots. In recent years, techniques such as joint learning, which consider the strong correlation between intent and slots, have yielded outstanding results. Certain approaches to joint slot filling and intent detection involve the sharing of parameters (Liu and Lane, 2016; Wang et al., 2018; Zhang and Wang, 2016). The association between intent detection and slot filling can be modeled through unidirectional interaction or bidirectional-flow interaction (Qin et al., 2021c).
Unidirectional interaction approaches (Goo et al., 2018; Li et al., 2018; Qin et al., 2019) primarily focus on the flow from intent to slot. Gating mechanisms have been employed as specialized functions to guide slot filling (Goo et al., 2018; Li et al., 2018). Qin et al. (2019) proposed a token-level intent detection model to mitigate error propagation.
Bidirectional-flow interaction models (E et al., 2019; Zhang et al., 2019; Liu et al., 2019; Qin et al., 2021a) consider the mutual impact between intent detection and slot filling. E et al. (2019) enhanced intent detection and slot filling bidirectionally through iteration mechanisms.
More recently, Chen et al. (2022) introduced a Self-distillation Joint SLU model, leveraging multi-task learning. They also treated multiple intent detection as a weakly supervised challenge, employing multiple instance learning (MIL). Huang et al. (2022) proposed a chunk-level intent detection framework, including an auxiliary task to identify intent transition points within utterances, thereby enhancing the recognition of multiple intents. And Cheng et al. (2023) proposed a transformative network built on the Transformer model, designed to reduce the complexity of multi-intent detection in SLU through scope recognition and bidirectional interaction between results of slot filling and intent detection.
A.2 Open Source Large Language Model
Advancements in natural language processing (NLP) have been significantly driven by the development of open-source Large Language Models (LLMs) like Llama2(Touvron et al., 2023), VicunaPeng et al. (2023), and MistralJiang et al. (2023). These models, with their extensive parameterization and training on diverse datasets, have expanded the capabilities in various NLP tasks.
Llama2, a creation of Meta AI, offers a range of models with parameters varying from 7 billion to 70 billion. It excels particularly in dialogue-based applications, utilizing auto-regressive language optimized transformers and advanced techniques such as supervised fine-tuning and reinforcement learning. This approach enhances Llama2’s performance, especially in dialogue generation and understandingTouvron et al. (2023). Vicuna, another notable open-source LLM, is specially designed for chatbot applications. It is trained on datasets that mirror real-world interactions, such as those from ShareGPT.com. Vicuna stands out for its transparent approach to training and its architectural innovations, offering various versions like Vicuna-13B to cater to different application needsPeng et al. (2023). Mistral 7B, developed by Mistral AI, is geared towards both English language processing and coding tasks. This 7.3 billion parameter model is distinguished for its superior performance over Llama 2 13B and Llama 1 34B in multiple benchmarks. The efficiency of Mistral 7B can be attributed to its use of grouped-query attention, which balances resource efficiency with output qualityJiang et al. (2023).
Chain of Thought Wei et al. (2022) is yet another LLM gaining attention for its capabilities in language generation. Together, these models, including Llama2, Vicuna, Mistral, and Chain of Thought, contribute to a diverse landscape of LLMs. They not only enhance NLP tasks but also have the potential to revolutionize content creation and search engine utilization.

### v3

@@ 5 Further Exploration
This section centres around an in-depth exploration of the LLM’s potential in advancing multi-intent SLU. Emphasizing consistency, we confine our experimental studies to the use of the EN-LLM series without incorporating human-written sub-intent instruction.
5.1 Influence of Different Intent Number
Within the ambit of both the MixATIS and MixSNIPS datasets, LLMs have exhibited a substantial competitive prowess over the baseline models, spanning a diverse range of performance metrics. In order to dissect the subtleties of these performance divergences, a systematic classification was adopted on the basis of the number of intents present in the MixATIS and MixSNIPS multi-intent datasets, as duly outlined in Table 2 and Table 3. Nonetheless, for the convenience of comparison with prior studies such as (Huang et al., 2022; Yin et al., 2024), this section primarily places its focus on the MixATIS dataset, though analogous conclusions can be drawn from the MixSNIPS dataset.
It becomes apparent that increased numbers of intents correspondingly escalate the challenge for all models, resulting in overall decreased performance metrics. In the context of a single-intent setting, the EN-Vicuña model exhibits a performance superiority over the Uni-MIS model by achieving an 80.4% overall accuracy rate, reflecting an advancement of 1.8%. And the EN-Mistral model unravels a noteworthy performance by outshining the Uni-MIS baseline by 4.8% in two-intent scenarios, thereby securing an overall accuracy of 55.3%. This leading position continues even into the three-intent setting, with the EN-Mistral model reaching an overall accuracy of 31.5%, which marginally underperforms Uni-MIS by only 0.2%.
Drawing from such observations, it suggests that LLMs exhibit a substantial capacity for proficiently handling complexities associated with a small number of intents. However, as the quantity of intents escalates, most models begin to display regressing performance, potentially attributable to the amplified error propagation in LLM’s generation process as a result of its autoregressive architecture.
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
@@ 5.1 Influence of Different Intent Number
Figure 3: Performance comparison of models on MixATIS and MixSNIPS datasets. The figures depict Semantic Accuracy, Slot F1, and Intent Accuracy across different data ratios for EN-Mistral, EN-Vicuna, and EN-Llama-2 models.
5.2 Influence of Training Data Ratio
@@
As illustrated in Figure 3, we probes the effect of varying volumes of training data on model proficiency, with a focus on three critical performance indicators: Semantic Accuracy, Slot F1 Score, and Intent Accuracy. Through progressive modulation of training data ratios at intervals of 0.2, 0.4, 0.6, 0.8, and 1.0. There appears to be a consistent and positive correlation between the data ratio and the amelioration of performance metrics across both the MixATIS and MixSNIPS datasets. This trend substantiates the hypothesis that an increase in dataset size can significantly bolster the model’s training efficacy, thereby refining accuracy and broadening the model’s generalization abilities. Upon closer inspection, the MixATIS and MixSNIPS datasets indicate that the EN-Mistral model, starting at a less optimal performance with a data ratio of 0.2, demonstrates notable advancements as the dataset size increases.
6 Introduction of ESA and CSA Metrics
@@
To enhance the evaluation of LLMs in the domain of multi-intent SLU, we propose two novel metrics for a more nuanced analysis: Entity Slot Accuracy and Combined Semantic Accuracy.
Entity Slot Accuracy (ESA): ESA is a key metric in the evaluation of LLM performance for multi-intent SLU tasks. Mathematically, ESA is determined by the ratio of correctly identified entity slots, denoted as $E_{c}$, to the total number of slots forecasted, denoted as $E_{t}$. The quantification can be articulated by:
$ESA=\frac{|E_{c}|}{|E_{t}|}\times 100\%$ (4)
Here, ||$\cdot$|| ⋅ | signifies the cardinality or the count of elements in a set. This concise representation of ESA hence encapsulates the model’s precision in parsing the semantic features of an utterance.
Combined Semantic Accuracy (CSA): CSA aims to provide an encompassing evaluation of a model’s adeptness at understanding both the intent and entity slots in an utterance. Denoting Intent Accuracy as $I_{a}$ and Entity Slot Accuracy as $E_{s}$, CSA is then mathematically represented as:
$CSA=I_{a}\times E_{s}$ (5)
This equation essentially encapsulates the model’s holistic accuracy in correctly identifying both the overall intent and the specific slots within a given utterance.
Upon analyzing the performance of various LLMs on the LM-MixATIS and LM-MixSNIPS datasets, the data reveals certain performance trends. The EN-Mistral model outshines others in the MixATIS dataset with an ESA of 60.6%, and a CSA of 48.8%. In the MixSNIPS dataset, EN-Mistral model achieves the highest ESA at 84.2% and the highest CSA at 82.2%. These scores reflect the model’s proficiency in handling the complexity of the dataset’s queries.
Notably, ESA and CSA provide a nuanced performance assessment, particularly focusing on the value of entity slots, and excluding the ’O’ tag, typically denoting tokens not corresponding to any entity slot in the utterance. This stringent focus endorses a more thorough evaluation of an LLM’s capabilities in discerning and accurately predicting relevant utterance information.
@@
EN-Vicuna 53.3 43.5 79.1 77.2
EN-Llama-2 53.1 43.8 79.3 76.8
EN-Mistral 60.6 48.8 80.2 78.3
ENSI-Vicuna 56.3 45.3 82.9 80.6
ENSI-Llama-2 50.0 39.7 83.9 81.9
ENSI-Mistral 53.1 41.5 84.2 82.2
@@
7 Related Work
7.1 Intent Detection and Slot Filling
The inherent interconnectedness of intent detection and slot filling has spurred the development of unified models that foster mutual interaction between the two elements. Joint learning techniques, acknowledging the potent correlation between intents and slots, have proven particularly efficacious in recent years. Certain methodologies facilitating simultaneous slot filling and intent detection employ shared parameters (Liu and Lane, 2016; Wang et al., 2018; Zhang and Wang, 2016), while others model the relationship between the two via either unidirectional interaction or bidirectional-flow interaction (Qin et al., 2021c).
Models adopting unidirectional interaction, such as those by (Goo et al., 2018; Li et al., 2018; Qin et al., 2019), primarily emphasize the flow from intent to slot. Gating mechanisms, functioning as specialized guiding forces for slot filling, have seen extensive use (Goo et al., 2018; Li et al., 2018). Qin et al. (2019) put forth a token-level intent detection model to curtail error propagation.
Bidirectional-flow interaction models (E et al., 2019; Zhang et al., 2019; Liu et al., 2019; Qin et al., 2021a), on the other hand, examine the reciprocal influence of intent detection and slot filling. E et al. (2019) utilized iterative mechanisms to enhance intent detection and slot filling in both directions.
Fine-grained intent detection and intent-slot interaction models have also seen remarkable advancements. Chen et al. (2022) developed a Self-distillation Joint SLU model exploitating multi-task learning, and treated multiple intent detection as a weakly-supervised problem solved through Multiple Instance Learning (MIL). Similarly, Huang et al. (2022) introduced a chunk-level intent detection framework that employs an auxiliary task to pinpoint intent transition points within utterances, thereby augmenting the recognition of multiple intents. Furthermore, Cheng et al. (2023) proposed a transformative network rooted in the Transformer model, designed to diminish the complexity of multi-intent detection in SLU. Recently, Yin et al. (2024) further develop an united multi-view intent-slot interaction framework(Uni-MIS), archiving promising performance.
7.2 Open Source Large Language Model
The landscape of Natural Language Processing (NLP) has been notably transformed by advancements in open-source Large Language Models (LLMs) such as Llama2(Touvron et al., 2023), VicunaPeng et al. (2023), and MistralJiang et al. (2023). These potent models, undergirded by vast parameterizations and trained on varied data corpuses, have significantly broadened the scope of NLP applications.
Llama2, pioneered by Meta AI, provides a spectrum of options, with models ranging from 7 to 70 billion parameters. Renowned for its prowess in dialog-centric tasks, Llama2 leverages auto-regressive transformers fine-tuned with supervised learning and bolstered by reinforcement learning techniquesTouvron et al. (2023). Vicuna sets its mark as a specialized open-source LLM for chatbot applications, honed through datasets reflective of authentic conversational exchanges—flagships of platforms like ShareGPT. Vicuna’s clear training protocols and unique architecture, with models such as Vicuna-13B, tailor its performance to specific conversational needsPeng et al. (2023). In contrast, Mistral AI’s Mistral 7B stands as a versatile model catering to both English language processing and coding assignments. This model, with 7.3 billion parameters, outshines peers such as Llama 2 13B and Llama 1 34B across diverse benchmarks, thanks to its innovative grouped-query attention mechanism that optimizes the balance between resource economy and model reliabilityJiang et al. (2023).
Accompanying these is the "Chain of Thought" method Wei et al. (2022), another vanguard in language generation. Collectively, models like Llama2, Vicuna, Mistral, and Chain of Thought fortify a varied array of LLMs. Their contribution extends beyond improving NLP tasks; they are poised to pioneer new horizons in content generation and redefine our engagement with search technologies.
@@ 8 Conclusion
In this investigation, we initially incorporated LLMs into the realm of multi-intent SLU, thus tailoring the LM-MixATIS and LM-MixSNIPS datasets to effectively align with the proficiencies of comprehensive models. Concurrently, we introduced two unique model series: EN-LLM and ENSI-LLM. The results of our research indicate that LLMs yield competitive outcomes when juxtaposed with robust baseline models prevalent in this field. Upon further examination of the impact of data rate and intent quantity, we discovered certain limitations associated with LLMs. In addition to these findings, we formulated evaluation metrics, specifically ESA and CSA, which prove crucial in accurately assessing the performance of LLMs.
@@ 9 Limitations
(1) Impact of LoRA on Performance: While leveraging Low-Rank Adaptation (LoRA) affords improved training efficiency and reduced memory usage, it does not always achieve the full performance potential afforded by tuning all model parameters. LoRA strictly focuses on the most impactful parameters, which, though efficient, can sometimes lead to a loss of subtle linguistic detail that comprehensive parameter tuning might capture. (2) Prospects for Improvement through Data Curation and Prompt Optimization: Our current research framework does not extend to the advanced strategies of selective data curation or intricate prompt engineering. Recognizing this as a limitation, we propose that future investigations will embrace these crucial techniques.
Appendix A Appendix
Model Slot(F1) Intent(Acc) Overall(Acc)
EN-ChatGPT (5-shot) 64.0 54.1 14.6
EN-Vicuna-13B 87.9 83.6 50.8
EN-Llama-2-13B 87.9 81.0 49.5
EN-Mistral-8*7B 87.3 80.7 48.4
Table 5: Performance on the MixATIS dataset using various Large Language Models.
Model Slot(F1) Intent(Acc) Overall(Acc)
EN-ChatGPT(5-shot) 62.9 83.9 12.7
EN-Vicuna-13B 95.9 97.5 80.7
EN-Llama-2-13B 96.7 97.8 83.3
EN-Mistral-8*7B 96.1 97.5 80.9
Table 6: Performance on the MixSNIPS dataset using various Large Language Models.
Note1: We also extend our evaluation to examine ChatGPT’s proficiency in a 5-shot learning context, paralleling the tests conducted with the LLMs outlined above. A distinctive aspect of our methodology with ChatGPT involves introducing an explicit intent and slot list to direct the model’s learning, enhancing its capability to align with specified NLP tasks.
Note2: Performance analysis across different scales of Large Language Models (LLMs) yields a compelling insight: within the MixATIS dataset, the transition to more substantial 13B and 8*7B models does not uniformly translate into significant gains over their 7B counterparts. This finding challenges the notion that escalating model size inherently overcomes the challenges in discerning slots and intents within single-domain datasets. Conversely, the multi-domain MixSNIPS dataset paints a different picture, where enhanced model sizes seemingly marshal improved outcomes. Notably, the EN-Vicuna-13B model marks a 1.8% increase in overall accuracy relative to its 7B counterpart, the EN-Llama-2-13B model outperforms by an impressive 4.4%, and the EN-Mistral-8*7B model uplifts its performance by 1.1% in overall accuracy. These contrasts suggest that larger models may have a distinctive edge in more straightforward, multi-domain settings, chiefly attributed to their superior aptitude for capturing and interpreting the complex mosaic of patterns within such varied datasets.